At a production process, the produced items are tested for defects. A defective unit is classified as such with probability 0.9, whereas a correct unit is classified as such with probability 0.85. Furthermore, 10% of the produced units are defective. What is the conditional probability that a unit is correct, given that is has been classified as defective?


1
Expert's answer
2021-12-17T13:25:33-0500

Let "A" be the event that an item is defective(correct) and "B" be the event that an item is classified defective.

The conditional probability that we need to find is "p(A|B)".

We are given that,

"p(A)=0.1\implies p(A')=1-p(A)=1-0.1=0.9"

"p(B|A)=0.9" and "p(B|A')=1-P(B'|A')=1-0.85=0.15"

To find "p(B)", we use the law of total probability as follows.

"P(B) = p(A)\times p(B|A) + p(A')\times p(B|A')=(0.1\times 0.9)+(0.9\times0.15)=0.225"

Since "p(A)\times p(B|A) = p(B)\times p(A|B)", we substitute for the values we have obtained above to get "p(A|B)".

Now,

"p(A|B)={p(A)\times p(B|A)\over p(B)}={0.1\times0.9\over 0.225}=0.4"

Therefore, the conditional probability that a unit is defective(correct), given that it has been classified as defective is 0.4.